# The inherent goodness of well educated intelligence

### 14 Jan 2024 v2

@@ I Introduction
There is an alternative – intelligence that “recognizes the global consequences of local actions”. General guidance is given by the leadership in the form of the Golden Rule. This rule recognizes that “the collective acts as one”, that one member’s costs are another member’s revenue, and vice versa. Zero sum, win-lose, local interactions will lead to a global lose-lose. There are only win-win or lose-lose global outcomes. The problem with this conservative (viscosity-less) control is that the equilibriums of maximum activity are unstable, like an inverted pendulum, and interactions of the collective with external system will lead to a relaxation to minimum activity. Leadership stimulates the economy of the collective to the meta stable maximum, but watches helplessly as the economy crashes into economic depressions.
Collective systems are not limited to social and economic systems. They also occur in physical, chemical, and biological systems. They are what are commonly referred to as complex systems with emergent or collective behaviors. The most well known physical collectives are the systems of elementary particles, often referred to Quantum Field Theories. There are also the classical physical collectives of interacting particles such as ideal gases and plasmas. These collectives can have $10^{25}$ or more members. This field of physics is referred to as Physical Kinetics or Plasma Physics. In order to have stable solutions, diffusive approximations or viscosity is added to the system. It is well known that this sacrifices physics for stability.
@@
This paper will first describe collective system as twinkling textures governed by a puppet master in Sec. II. How to control these textures will be presented in Sec. III, and the underlying analytic theory shown in Sec. IV. A more detailed history of social and economic collectives will be given in Sec. V, followed by a detailed discussion of the ethos of stable and unstable equilibriums in Sec. VI. There will be a general discussion with conclusions in Sec. VII
@@ II A collection of twinkling textures
Evolution of a collective system made up of a large number of identical individual members has, as a foundation, the symmetries of the individual. This is displayed in Fig. 1 where the prototype individual or puppet master moves according to a flow generated by the symmetry of the motion with corresponding constants of the motion. External forces push on the puppet master in a direction perpendicular to the motion. There will be equal number of stable and unstable points of equilibrium where the puppet master will effectively have infinite mass and the external force will not be able to move the puppet master. Although the external force can change the constants of the motion, it can not change the points of equilibrium leading to true invariants of the motion. The puppet master will be characterized by two types of strings that it can pull: co-state strings $p_{i}$, and state strings $q_{i}$. These strings are transformed into fundamental co-states $P_{i}$ and states $Q_{i}$, where the $Q$’s are the direction of the motion of the puppet master and the $P$’s are the direction that the external force can push the puppet master. This flow is generated by a function $S_{P}(q)$ determined by the Hamilton-Jacobi-Bellman (HJB) equation (Goldstein, 1980). Mathematical details will be discussed in Sec. IV.
@@
Figure 4: Topography of geodesic motion of an electron about an ion in a constant magnetic field. (a) motion shown as the red lines with the black arrows, separatrix shown as the thick black line with black arrows, stable equilibriums shown as white o-points, meta stable equilibriums shown as white x-points, external thermal force shown as yellow arrows. (b) illustration showing choice at the crossroads, take the easy road to the valley center or the hard road to the mountain pass. (c) path shown in purple with thermal force that reaches thermal equilibrium near the valley center.
@@ III Control of twinkling textures
There are three approaches to controlling conservative systems: viscous, feedback, and ponderomotive control. Starting in the 14th and 15th centuries in Renaissance Northern Italy, lead by the Strozzis and Medicis, viscous economic control became popular. This is more commonly known as usury or borrow/loan financing. It is like controlling a car by going full gas, then modulating the brakes to control the speed of the car. It is obviously very inefficient. Financially, it puts a free cash flow constraint on the operations of the economies that most times leads to a 90% to 95% reduction in sustainable economic activity. What is even more grave is a dissipation of revenue that is otherwise needed for sustenance and growth of the economy. This is a self fulfilling prophesy that leads to a premature demise of the economy. An example of this degradation in performance, for a producer of unconventional hydrocarbons, is given in App. A of Glinsky and Sievert (2023).
Economic and social viscosity takes many forms. It is not only usury (that is debt), it is servitude, slavery, intolerance, and lack of freedom. Society is locked up, not only figuratively, but often literally. It is also used to control instabilities in the numerical solution of partial differential equations (PDEs), but is well know to sacrifice physics for stability. There is also a whole industry that uses viscosity to solve the HJB equation (Bardi et al., 1997). Viscosity lowers the performance of both x-points and o-points. If the viscosity is greater than a critical amount, it removes the finite x-points and o-points from the system.
@@ IV The underlying analytic theory
There are two parts to the theory in Fig. 1. The first is the logarithmic or canonical generating functional $S_{p}[f(x)]$ called the Heisenberg Scattering Transformation (HST). This functional deeply deconvolves the twinkling textures of the collective $[\pi_{i}(x),f_{i}(x)]$ to the canonical coordinates $(p_{i},q_{i})$ of the puppet master. It also can be viewed as imaging all the multiple reflections of the puppet master seen through the hall of mirrors back to the primary image of the puppet master. In the language of Kinetic Theory or Plasma Physics, this is the Mayer Cluster Expansion in terms of the $m$-body correlations. In the language of Quantum Field Theory, this is Heisenberg’s S-Matrix ($m$-body Scattering Cross Sections or $m$-body Green’s Function), or the Wigner-Weyl Transformation. The formula for the HST is
@@
where $\phi$ is a Father Wavelet or window function, $\psi_{p}$ is a scaled Mother Wavelet forming an orthogonal set of coherent states, the convolution is given by
@@
With the definition
$R_{0}(z)\equiv z+\text{e}^{\text{i}\arg(z)},$ (5)
@@
Because of the logarithmic structure of $S_{p}[f(x)]$ the canonical motion will be confined to a complex linear hyperplane $\mathbb{C}^{n}$, where $n$ is the number of fields, with basis vectors $\left|\beta_{i}(z)\right>$ that can be found with a Principal Components Analysis (PCA). These basis vectors can be viewed as the solutions to the Renormalization Group Equations (RGE) (Weinberg, 2005) or as the Singularity Spectrums. This logarithmic generating functional can also be viewed as the logarithmic field distribution function
@@
as is done by AI generators like GANs and GPTs. The HST is shown diagrammatically as a Mayer Cluster Expansion in Fig. 6. It should be noted that $|z|$ is a two sided Rectified Linear Unit (ReLU), so that the HST has the form of a Convolutional Neural Network where the activation function is $\text{i}\,\ln R_{0}$, and the pooling operator is $\phi\star$.
@@
$S(P,E;q,\tau)\equiv S_{P}(q)-E(P)\tau,$ (10)
and $W(P;q)\equiv S_{P}(q)$. The corresponding Poincaré one form is $\lambda\equiv p\,dq-E\,d\tau$ on extended phase space $\mathscr{H}=H\otimes\text{Ad}(H)$. The generating function $S_{P}(q)$ can be identified as the action, entropy and
@@
associated with the complex Lie group $\mathscr{H}=H\otimes\text{Ad}(H)$. The HJB equation is very difficult to solve analytically, but it is very easy to approximate the solution with a Multi-Layer Perceptron (MLP) (Goodfellow et al., 2016) with ReLU activation. The MLP/ReLU is a piecewise linear function approximator that is well suited to approximating an analytic function that is maximally flat, a solution to Laplace’s equation, except for a few isolated singularities that will have a discontinuity in the first derivative at those singularities. A block diagram of how to solve for the solution of the HJB equation using a several MLPs, which is effectively a decoder into a $(P,E;Q,\tau)$ Reduced Order Model (ROM), followed by the propagator, then a matched encoder to $(p,q)$, is shown in Fig. 7.
@@ V History of social and economic collectives
As is shown in Fig. 8, US society was freed in the 1960’s. Jimmy Carter was a very well intentioned shepard when he was elected president in 1976. Unfortunately, he had no tools to control such a frictionless society and watched it wonder out of control with large amounts of inflation. The response of the political right was to reimpose large amounts of friction on US society. The implementation plan we put forward in the 1971 manifesto of Lewis Powell (Andersen, 2021), put into action by the cabal of the Koch brothers, Olin and Coors, rolled out with the election of Ronald Reagan in 1980, and is reaching its natural endpoint with the fascism of Donald Trump.

### 19 Jan 2024 v3

@@ I Introduction
There is an alternative – intelligence that “recognizes the global consequences of local actions”. General guidance is given by the leadership in the form of the Golden Rule. This rule recognizes that “the collective acts as one”, that one member’s costs are another member’s revenue, and vice versa. Zero sum, win-lose, local interactions will lead to a global lose-lose. There are only win-win or lose-lose global outcomes. The problem with this conservative (viscosity-less) control is that the equilibriums of maximum activity are unstable, like an inverted pendulum, and interactions of the collective with external systems will lead to a relaxation to minimum activity. Leadership stimulates the economy of the collective to the metastable maximum, but watches helplessly as the economy crashes into economic depressions.
Collective systems are not limited to social and economic systems. They also occur in physical, chemical, and biological systems. They are what are commonly referred to as complex systems with emergent or collective behaviors. The most well known physical collectives are the systems of elementary particles, often referred to Quantum Field Theories (Weinberg, 2005). There are also the classical physical collectives of interacting particles such as ideal gases and plasmas. These collectives can have $10^{25}$ or more members. This field of physics is referred to as Physical Kinetics (Lifschitz and Pitajewski, 1983) or Plasma Physics (Nicholson, 1983). In order to have stable solutions, diffusive approximations or viscosity is added to the system. It is well known that this sacrifices physics for stability.
@@
This paper will first describe a collective system as twinkling textures governed by a puppet master in Sec. II. How to control these textures will be presented in Sec. III, and the underlying analytic theory shown in Sec. IV. A more detailed history of social and economic collectives will be given in Sec. V, followed by a detailed discussion of the ethos of stable and unstable equilibriums in Sec. VI. There will be a general discussion with conclusions in Sec. VII. As an appendix, there is a discussion of the Simple Beauty of the Crazy Ones (App. A).
@@ II A collection of twinkling textures
Evolution of a collective system made up of a large number of identical individual members has, as a foundation, the symmetries of the individual. This is displayed in Fig. 1 where the prototype individual or puppet master moves according to a flow generated by the symmetry of the motion with corresponding constants of the motion. External forces push on the puppet master in a direction perpendicular to the motion. There will be equal number of stable and unstable points of equilibrium where the puppet master will effectively have infinite mass and the external force will not be able to move the puppet master. Although the external force can change the constants of the motion, it can not change the points of equilibrium leading to true invariants of the motion. The puppet master will be characterized by two types of strings that it can pull: co-state strings $p_{i}$, and state strings $q_{i}$. These strings are transformed into fundamental co-states $P_{i}$ and states $Q_{i}$, where the $Q$’s are the direction of the motion of the puppet master and the $P$’s are the direction that the external force can push the puppet master. This flow is generated by a function $S_{P}(q)$ determined by the Hamilton-Jacobi-Bellman (HJB) equation (Goldstein, 1980; Lichtenberg and Lieberman, 2010). Mathematical details will be discussed in Sec. IV.
@@
Figure 4: Topography of geodesic motion of an electron about an ion in a constant magnetic field: (a) motion shown as the red lines with the black arrows, separatrix shown as the thick black line with black arrows, stable equilibriums shown as white o-points, metastable equilibriums shown as white x-points, external thermal force shown as yellow arrows, (b) illustration showing choice at the crossroads, take the easy road to the valley center or the hard road to the mountain pass, (c) path shown in purple with thermal force that reaches thermal equilibrium near the valley center.
@@ III Control of twinkling textures
There are three approaches to controlling conservative systems: viscous, feedback, and ponderomotive control. Starting in the 14th and 15th centuries in Renaissance Northern Italy, lead by the Strozzis and Medicis, viscous economic control became popular. This is more commonly known as usury or borrow/loan financing. It is like controlling a car by going full gas, then modulating the brakes to control the speed of the car. It is obviously very inefficient. Financially, it puts a free cash flow constraint on the operations of the economies that most times leads to a 90% to 95% reduction in sustainable economic activity. What is even more grave is a dissipation of revenue that is otherwise needed for sustenance and growth of the economy. This is a self fulfilling prophesy that leads to a premature demise of the economy, as shown in Eq. 15. An example of this degradation in performance, for a producer of unconventional hydrocarbons, is given in App. A of Glinsky and Sievert (2023).
Economic and social viscosity takes many forms. It is not only usury (that is debt), it is servitude, slavery, intolerance, and lack of freedom. Society is locked up, not only figuratively, but often literally. It is also used to control instabilities in the numerical solution of partial differential equations (PDEs), but is well known to sacrifice physics for stability. There is also a whole industry that uses viscosity to solve the HJB equation (Bardi et al., 1997). Viscosity lowers the performance of both x-points and o-points. If the viscosity is greater than a critical amount, it removes the finite x-points and o-points from the system.
@@ IV The underlying analytic theory
There are two parts to the theory in Fig. 1. The first is the logarithmic or canonical generating functional $S_{p}[f(x)]$ called the Heisenberg Scattering Transformation (HST). This functional deeply deconvolves the twinkling textures of the collective $[\pi_{i}(x),f_{i}(x)]$ to the canonical coordinates $(p_{i},q_{i})$ of the puppet master. It also can be viewed as imaging all the multiple reflections of the puppet master seen through the hall of mirrors back to the primary image of the puppet master. In the language of Kinetic Theory or Plasma Physics, this is the Mayer Cluster Expansion (Uhlenbeck and Ford, 1963) in terms of the $m$-body correlations. In the language of Quantum Field Theory, this is Heisenberg’s S-Matrix (Chew, 1961; Weinberg, 2005) ($m$-body Scattering Cross Sections or $m$-body Green’s Functions), or the Wigner-Weyl Transformation (Wigner, 1932; Weyl, 1950). The formula for the HST is
@@
where $\phi$ is a Father Wavelet or window function, $\psi_{p}$ is a scaled Mother Wavelet forming an orthogonal set of coherent states (Ali et al., 2000; Mallat, 1999), the convolution is given by
@@
With the definition of $R_{0}$ given in Eq. 3,
@@
Because of the logarithmic structure of $S_{p}[f(x)]$ the canonical motion will be confined to a complex linear hyperplane $\mathbb{C}^{n}$, where $n$ is the number of fields, with basis vectors $\left|\beta_{i}(z)\right>$ that can be found with a Principal Components Analysis (PCA). These basis vectors can be viewed as the solutions to the Renormalization Group Equations (RGE) (Weinberg, 2005) or as the Singularity Spectrums (Mallat, 1999). This logarithmic generating functional can also be viewed as the logarithmic field distribution function
@@
as is done by AI generators like GANs and GPTs. The HST is shown diagrammatically as a Mayer Cluster Expansion in Fig. 6. It should be noted that $|z|$ is a two sided Rectified Linear Unit (ReLU), so that the HST has the form of a Convolutional Neural Network (CNN) where the activation function is $\text{i}\,\ln R_{0}$, and the pooling operator is $\phi\star$.
@@
$S(P,E;q,\tau)\equiv S_{P}(q)-E(P)\tau.$ (9)
The corresponding Poincaré one form is $\lambda\equiv p\,dq-E\,d\tau$ on extended phase space $\mathscr{H}=H\otimes\text{Ad}(H)$. The generating function $S_{P}(q)$ can be identified as the action, entropy or
@@
associated with the complex Lie group $\mathscr{H}=H\otimes\text{Ad}(H)$. The HJB equation is very difficult to solve analytically, but it is very easy to approximate the solution with a Multi-Layer Perceptron (MLP) (Goodfellow et al., 2016) with ReLU activation. The MLP/ReLU is a piecewise linear universal function approximator that is well suited to approximating an analytic function that is maximally flat, a solution to Laplace’s equation, except for a few isolated singularities that will have a discontinuity in the first derivative at those singularities. A block diagram of how to solve for the solution of the HJB equation using several MLPs, which is effectively a decoder into a $(P,E;Q,\tau)$ Reduced Order Model (ROM), followed by the propagator, then a matched encoder to $(p,q)$, is shown in Fig. 7.
@@ V History of social and economic collectives
As is shown in Fig. 8, US society was freed in the 1960’s. Jimmy Carter was a very well intentioned shepard when he was elected president in 1976. Unfortunately, he had no tools to control such a friction-less society and watched it wonder out of control with large amounts of inflation. The response of the political right was to reimpose large amounts of friction on US society. The implementation plan was put forward in the 1971 manifesto of Lewis Powell (Andersen, 2021), put into action by the cabal of the Koch brothers, Olin and Coors, rolled out with the election of Ronald Reagan in 1980, and is reaching its natural endpoint with the fascism of Donald Trump.
@@ Acknowledgements.
Appendix A Simple Beauty and the Crazy Ones
The correct theory is the “simply beautiful” theory as postulated by Physics Nobel Laureates Murray Gell-Mann and Paul Dirac. Simple can be Frank Lloyd Wright’s Fallingwater, Phillip Johnson’s postmodern Pennzoil building, Steve Jobs’ electronic devices, a Danish Bang & Olufsen stereo, a Piet Mondrian painting, or self organized fluid flow constrained by topological invariants (see Fig. 9). There is also the simple beauty of textures, that is Benoit Mandelbrot’s fractals, a Jackson Pollock painting, a Brazilian favela, the chaos boards of Carrie Mathison in “Homeland” or John Nash in “A Beautiful Mind”, or Andrey Kolmogorov’s well developed turbulence (see Fig. 10). The later is as simple as the former. There is a high degree of organization or correlation in the later. Our visual and audio cortex, along with the Convolutional Neural Networks (CNNs) of modern AI are very good at identifying the simple beauty or symmetry in these complex images or signals.
Figure 9: Examples of simple geometric structures: (a) Frank Lloyd Wright’s Fallingwater, (b) Phillip Johnson’s postmodern Pennzoil building, (c) Steve Jobs’ iMac, (d) a Bang & Olufsen stereo, (e) a Piet Mondrian painting, and (f) self organized fluid structures constrained by vorticity and helicity.
Figure 10: Examples of simple textures: (a) a Brazilian favela, (b) a Jackson Pollock painting, (c) a fractal pattern, (d) well developed turbulent fluid flow, (e) chaos boards of Carrie Mathison and John Nash, and (f) the interior and exterior of Antoni Gaudí’s La Sagrada Familia.
The beautiful simplicity of these textures is not always recognized. My great grandmother was committed to an insane asylum in western Pennsylvania at 40 years of age where she remained for the rest of her life, because of the simple rambling textures of her speech, which the medical professionals found incoherent and the manifestation of insanity – she was crazy. She was from Transylvania and speaking in Romanian. Are the hallucinations of schizophrenia, and the multiple personalities of Dissociative Identity Disorder just alternative textures of a high functioning creative mind? How often do we confuse genius with insanity? Is insanity, really genius?
For complex or conservative collective systems there are two types of simply beautiful emergent behaviors. The first comes from an inverse cascade of energy to larger scale and a relaxation of the system to self organized structures. These are the local minimums, stable equilibriums, or o-points – the simple geometric structures shown in Fig. 9. The second comes from a normal cascade of energy to smaller scale and a relaxation of the system to a turbulent structure. These are the local maximums, metastable equilibriums, or x-points – the simple textures shown in Fig. 10. Both are exposing the homology classes of the topology $\beta^{*}$. The simple beauty is the symmetry or topology of the system quantified by the constants and their analytic continuation $H(\beta)$, the homology classes $\beta^{*}$, or the generators of the canonical flow and the constants $S_{p}[f(x)]$ and $H(p,q)$ or $S_{P}(q)$ and $E(P)$.
These are all equivalent ways of quantifying the symmetry, that is the topology of the system. Given the boundary $\beta^{*}$ where $\omega^{*}_{Q}=0$, that is the singular points or equilibriums, $H(\beta)$ can be found by solving the Laplace’s equation. Given the constants $H(p,q)=E(P)$, that is $\text{Re}(H(\beta))$, $\text{Im}(H(\beta))=Q=\partial S/\partial P$ can be found by analytic continuation. Given the constants $H(p,q)$, the action $S_{P}(q)$ can be found by solving the HJB equation. Remember that the action is also called the propagator, the generator of the canonical flow, the entropy, or the logarithm of the density.